# AT 2025abao: the fourth luminous red nova in M 31

### v2

@@ Acknowledgements.
AR acknowledges financial support from the GRAWITA Large Program Grant (PI D’Avanzo). AR, AP, GV, NER acknowledge financial support from the PRIN-INAF 2022 ”Shedding light on the nature of gap transients: from the observations to the models”. NER acknowledges support from the Spanish Ministerio de Ciencia e Innovación and the Agencia Estatal de Investigación 10.13039/501100011033 under the program Unidad de Excelencia María de Maeztu CEX2020-001058-M. TMR acknowledges support from the Research Council of Finland project 350458 and the Cosmic Dawn Center, which is funded by the Danish National Research Foundation under grant DNRF140. MDS is funded by the Independent Research Fund Denmark (grant nr. 10.46540/2032-00022B). YZC is supported by the National Natural Science Foundation of China (NSFC, Grant No. 12303054), the National Key Research and Development Program of China (Grant No. 2024YFA1611603), the Yunnan Fundamental Research Projects (Grant Nos. 202401AU070063, 202501AS070078), and the International Centre of Supernovae, Yunnan Key Laboratory (No. 202302AN360001). Based on observations collected at Copernico and Schmidt telescopes (Asiago Mount Ekar, Italy) of the INAF – Osservatorio Astronomico di Padova, and at the Galileo telescope of the Padova Univesity (Asiago Mount Pennar, Italy). AR thanks Alessandro Fabris and Virginia Albanese for having participated to the observations with AFOSC. Support to ATLAS was provided by NASA grant NN12AR55G. Based on observations made with the NOT, owned and operated jointly by Aarhus University, the University of Turku, the University of Oslo, the University of Iceland and Stockholm University, under program 72-305 (PI Reguitti), and via the NUTS2 collaboration which is supported in part by the Instrument Centre for Danish Astrophysics, and the Finnish Centre for Astronomy with ESO via Academy of Finland grant nr. 306531. Based on observations made with the GTC. Based in part on observations made at the TNG, operated by INAF, under program A50TAC_41 (PI Valerin). All three telescopes are installed at the Spanish Observatorio del Roque de los Muchachos of the Instituto de Astrofísica de Canarias, on the island of La Palma. We acknowledge the use of data from the Swift data archive. FR is grateful to iTelescope.Net for providing observing time to use their remote telescopes and to the AAVSO who made it possible to use the AAVSOnet telescopes. SPHEREx is a joint project of the Jet Propulsion Laboratory and the California Institute of Technology, and is funded by the National Aeronautics and Space Administration.References
@@
- The SPHEREx Image and Spectrophotometry Processing Pipeline. arXiv e-prints, pp. arXiv:2511.15823. Note: Submitted to ApJS External Links: Document, 2511.15823, ADS entry Cited by: §3.7.
- Evidence for the merger hypothesis in V4332 Sgr: a low 12C/13C ratio and multiple outbursts. arXiv e-prints, pp. arXiv:2602.14781. External Links: 2602.14781, ADS entry Cited by: §3.2.2.
@@
- The luminous red nova AT 2018bwo in NGC 45 and its binary yellow supergiant progenitor. A&A 653, pp. A134. External Links: Document, 2102.05662, ADS entry Cited by: §1, §3.4, §3.5, §3.8, §5, §5.
@@
- The SPHEREx Satellite Mission. arXiv e-prints, pp. arXiv:2511.02985. Note: accepted by ApJ External Links: Document, 2511.02985, ADS entry Cited by: §3.7.
@@
- Spectral response of SPHEREx. arXiv e-prints, pp. arXiv:2602.09139. External Links: 2602.09139, ADS entry Cited by: §3.7.
@@
- Hot springs and dust reservoirs: JWST reveals the dusty, molecular aftermath of extragalactic stellar mergers. arXiv e-prints, pp. arXiv:2508.03932. Note: Submitted to ApJ External Links: Document, 2508.03932, ADS entry Cited by: §5.
@@
- The Slow Brightening of WNTR23bzdiq/WTP19aalzlk: Possible Onset of Common Envelope Evolution in an Asymptotic Giant Branch Star?. ApJ 993 (1), pp. 109. External Links: Document, 2505.09691, ADS entry Cited by: Appendix B, §1, §2, Figure 2, §3.1, §3.2.1, §3.4, §3.7, Figure 15, §5, §5, §5, §5, §5.
@@
- 0.8-13 Micron Spectroscopy of V838 Monocerotis and a Model for Its Emission. ApJ 607 (1), pp. 460–473. External Links: Document, ADS entry Cited by: §3.8.
@@
- Shock-powered light curves of luminous red novae as signatures of pre-dynamical mass-loss in stellar mergers. MNRAS 471 (3), pp. 3200–3211. External Links: Document, 1705.03895, ADS entry Cited by: §3.2.2.
@@
- Forbidden hugs in pandemic times. IV. Panchromatic evolution of three luminous red novae. A&A 671, pp. A158. External Links: Document, 2208.02782, ADS entry Cited by: §3.4, §3.5, §3.8, §5, footnote 11.
@@
- The Infrared Telescope Facility (IRTF) Spectral Library: Cool Stars. ApJS 185 (2), pp. 289–432. External Links: Document, 0909.0818, ADS entry Cited by: §3.7.
@@
- Spectral Synthesis of TiO Lines. ApJ 498 (2), pp. 851–862. External Links: Document, ADS entry Cited by: §3.8.
@@ Appendix A Lines in the Na i D region in the Echelle spectra
In the first Echelle spectrum of AT 2025abao at d (Figure 18, top panel), a prominent absorption feature with a profile and intensity similar to the Na i D 2 line due to the transient is visible at 5890.9 Å (in the host galaxy rest frame). This has been tentatively identified in Figure 18 (top panel) as the transition Na i D 2 from a third Na I doublet, redshifted by 50 km s-1 with respect to that of the LRN. However, the corresponding Na i D 1 transition, which would be located at 5896.9 Å, is not detected. Therefore, we believe that this feature is not an additional Na i D component, but most likely an absorption metal line. One possibility (though uncertain) is Fe ii 5891.48 (multiplet 211), which is consistent with the detection of other Fe ii lines in the spectrum. However, no other Fe ii lines from the same ionization stage has been found.
In the second Echelle spectrum at d (Figure 18, bottom panel), the 5890.9 line is still present, but a new absorption is now visible at 5896.9 Å, which is fortuitously at the expected position of the Na i D 1 transition assuming that the doublet is redshifted by 50 km s-1 with respect to the transient. However, accounting that more absorption lines of metals in neutral state have appeared in this phase, as the continuum temperature decreases, we tentatively identify this new line as Fe i 5896.74. Though an identification as V ii 5897.54 is also possible, as absorption bands from the VO molecule are observed in the spectra at later phases.
@@ Appendix B Updated progenitor mass – luminosity relation
Considering the progenitor of AT 2025abao identified in archival images before the outburst, and its mass determined independently by K25, we can add it to the sample of objects used by Cai et al. (2022a) to derive the empirical relation between the progenitor mass and -band absolute magnitude at the second peak (or during the plateau phase) for LRNe, with the goal of updated the with the robust data inferred for AT 2025abao. Adopting their estimate of , and our value of mag, the updated Mass-Luminosity relation becomes:

### v3

@@ Acknowledgements.
We thank the anonymous referee for the useful comments and suggestions to improve the manuscript. AR, GV, YZC acknowledge financial support from the SOXS project (PI S. Campana). AR, AP, GV, NER acknowledge financial support from the PRIN-INAF 2022 ”Shedding light on the nature of gap transients: from the observations to the models”. NER acknowledges support from the Spanish Ministerio de Ciencia e Innovación and the Agencia Estatal de Investigación 10.13039/501100011033 under the program Unidad de Excelencia María de Maeztu CEX2020-001058-M. TMR acknowledges support from the Research Council of Finland project 350458 and the Cosmic Dawn Center, funded by the Danish National Research Foundation under grant DNRF140. MDS is funded by the Independent Research Fund Denmark (grant nr. 10.46540/2032-00022B). YZC is supported by the National Natural Science Foundation of China (Grant N. 12303054), the National Key Research and Development Program of China (N. 2024YFA1611603), the Yunnan Fundamental Research Projects (N. 202501AS070078), and the International Centre of Supernovae, Yunnan Key Laboratory (N. 202302AN360001). Based on observations collected at Copernico and Schmidt telescopes (Asiago Ekar, Italy) of the INAF-Osservatorio Astronomico di Padova, and at the Galileo telescope of the Padova Univesity (Asiago Pennar, Italy). AR thanks Alessandro Fabris and Virginia Albanese for having participated to the observations with AFOSC. Support to ATLAS was provided by NASA grant NN12AR55G. Based on observations made with the NOT, owned and operated jointly by Aarhus University, Turku University, Oslo University, the University of Iceland and Stockholm University, under program 72-305 (PI Reguitti), and via the NUTS2 collaboration which is supported in part by the Instrument Centre for Danish Astrophysics, and the Finnish Centre for Astronomy with ESO via Academy of Finland grant nr. 306531. Based on observations made with the GTC. Based in part on observations made at the TNG, operated by INAF, under program A50TAC_41 (PI Valerin). All three telescopes are installed at the Spanish Observatorio del Roque de los Muchachos of the Instituto de Astrofísica de Canarias, on the island of La Palma. We acknowledge the use of data from the Swift data archive. FR is grateful to iTelescope.Net for providing observing time to use their remote telescopes and to the AAVSO who made it possible to use the AAVSOnet telescopes. SPHEREx is a joint project of the Jet Propulsion Laboratory and the California Institute of Technology, and is funded by the National Aeronautics and Space Administration.References
@@
- The SPHEREx Image and Spectrophotometry Processing Pipeline. arXiv e-prints, pp. arXiv:2511.15823. Note: Submitted to ApJS External Links: Document, 2511.15823, ADS entry Cited by: §3.7.1.
- The Birth of High-Mass Stars: Accretion and/or Mergers?. AJ 129 (5), pp. 2281–2293. External Links: Document, astro-ph/0502485, ADS entry Cited by: §5.
- Evidence for the Merger Hypothesis in V4332 Sgr: A Low 12C/13C Ratio and Multiple Outbursts. ApJ 999 (2), pp. L35. External Links: Document, 2602.14781, ADS entry Cited by: §3.2.2.
@@
- An Intermediate Luminosity Transient in NGC 300: The Eruption of a Dust-Enshrouded Massive Star. ApJ 699 (2), pp. 1850–1865. External Links: Document, 0901.0710, ADS entry Cited by: §4.1.
@@
- The luminous red nova AT 2018bwo in NGC 45 and its binary yellow supergiant progenitor. A&A 653, pp. A134. External Links: Document, 2102.05662, ADS entry Cited by: §1, §3.4, §3.5, §3.7.2, §5, §5.
@@
- The SPHEREx Satellite Mission. ApJ 999 (1), pp. 139. External Links: Document, 2511.02985, ADS entry Cited by: §3.7.1.
@@
- From Common Envelope Evolution to Luminous Red Novae. I. A One-dimensional Radiation Hydrodynamic Model. ApJ 999 (2), pp. 217. External Links: Document, ADS entry Cited by: §3.2.2.
@@
- Journey to the center of the common envelope evolution: Inner dynamics of the post-dynamical inspiral. A&A 697, pp. A68. External Links: Document, 2412.04419, ADS entry Cited by: §4.1.
@@
- Spectral response of SPHEREx. arXiv e-prints, pp. arXiv:2602.09139. Note: Submitted to ApJS External Links: 2602.09139, ADS entry Cited by: §3.7.1.
@@
- Hot Springs and Dust Reservoirs: JWST Reveals the Dusty, Molecular Aftermath of Extragalactic Stellar Mergers. ApJ 999 (1), pp. 16. External Links: Document, 2508.03932, ADS entry Cited by: §5.
@@
- The Slow Brightening of WNTR23bzdiq/WTP19aalzlk: Possible Onset of Common Envelope Evolution in an Asymptotic Giant Branch Star?. ApJ 993 (1), pp. 109. External Links: Document, 2505.09691, ADS entry Cited by: Appendix B, §1, §2, Figure 2, §3.1, §3.2.1, §3.4, §3.7.1, Figure 15, §5, §5, §5, §5, §5.
- Type II intermediate-luminosity optical transients (ILOTs). MNRAS 467 (3), pp. 3299–3305. External Links: Document, 1611.05855, ADS entry Cited by: §5.
@@
- 0.8-13 Micron Spectroscopy of V838 Monocerotis and a Model for Its Emission. ApJ 607 (1), pp. 460–473. External Links: Document, ADS entry Cited by: §3.7.2.
@@
- Revisiting the V1309 Sco 2008 outburst spectra. Observational evidence for the theoretical modeling of stellar mergers. A&A 664, pp. A12. External Links: Document, 2206.03191, ADS entry Cited by: §4.1.
@@
- Shock-powered light curves of luminous red novae as signatures of pre-dynamical mass-loss in stellar mergers. MNRAS 471 (3), pp. 3200–3211. External Links: Document, 1705.03895, ADS entry Cited by: §3.2.2, §4.1.
@@
- Forbidden hugs in pandemic times. IV. Panchromatic evolution of three luminous red novae. A&A 671, pp. A158. External Links: Document, 2208.02782, ADS entry Cited by: §3.4, §3.5, §3.7.2, §5, footnote 11.
@@
- The Infrared Telescope Facility (IRTF) Spectral Library: Cool Stars. ApJS 185 (2), pp. 289–432. External Links: Document, 0909.0818, ADS entry Cited by: §3.7.1.
@@
- Eruptions of the V838 Mon type: stellar merger versus nuclear outburst models. A&A 451 (1), pp. 223–236. External Links: Document, astro-ph/0509379, ADS entry Cited by: §5.
@@
- Spectral Synthesis of TiO Lines. ApJ 498 (2), pp. 851–862. External Links: Document, ADS entry Cited by: §3.7.2.
@@ Appendix A Lines in the Na i D region in the Echelle spectra
In the first Echelle spectrum of AT 2025abao at d (Fig. 18, top panel), a prominent absorption feature with a profile and intensity similar to the Na i D2 line due to the transient is visible at 5890.9 Å (in the host galaxy rest frame). This has been tentatively identified as the transition Na i D2 from a third Na I doublet, redshifted by 50 km s-1 with respect to that of the LRN. However, the corresponding Na i D1 transition, which would be located at 5896.9 Å, is not detected. Therefore, we believe that this feature is not an additional Na i D component, but most likely an absorption metal line. One possibility (though uncertain) is Fe ii 5891.48 (multiplet 211), which is consistent with the detection of other Fe ii lines in the spectrum. However, no other Fe ii lines from the same ionisation stage has been found.
In the second Echelle spectrum at d (Fig. 18, bottom panel), the 5890.9 line is still present, but a new absorption is now visible at 5896.9 Å, which is fortuitously at the expected position of the Na i D1 transition, assuming that the doublet is redshifted by 50 km s-1 with respect to the transient. However, accounting that more absorption lines of metals in neutral state had appeared at this phase, as the continuum temperature decreased, we tentatively identify this new line as Fe i 5896.74, though an identification as V ii 5897.54 is also possible, as absorption bands from the VO molecule are observed in the spectra at later phases.
@@ Appendix B Updated progenitor mass – luminosity relation
Considering the progenitor of AT 2025abao identified in archival images before the outburst, and its mass determined independently by K25, we can add it to the sample of objects used by Cai et al. (2022a) to derive the empirical relation between the progenitor mass and -band absolute magnitude at the second peak (or during the plateau phase) for LRNe, with the goal of updating it with the robust data inferred for AT 2025abao. Adopting their estimate of , and our value of mag, the updated mass-luminosity relation becomes: